OpenStudy (anonymous):

Hi there May I know how to integrate this, 1/((x^2-9)^(1/2)). please note, this is definite integral. x is, [2*sqrt(3)] to [3*sqrt(2)]

OpenStudy (mathteacher1729):

This looks like a trig-sub integral. Have you sketched a reference triangle?

OpenStudy (mathteacher1729):

Here is a summary of the trig subs I made for my students. :)

OpenStudy (angela210793):

=\[\int\limits_{2\sqrt{3}}^{3\sqrt{2}}1/\sqrt{x^2-9} dx\]

OpenStudy (anonymous):

you can use online wolfram alpha

OpenStudy (anonymous):

exactly I went to upto Integral of (sec ). But I cant work out after that step.

OpenStudy (angela210793):

how abt replacing sqr(x^2-9) with a t?

OpenStudy (mathteacher1729):

Angela if you let t = sqrt ( x - 9) dt does not equal a constant, so that won't work. You need a trig sub.

OpenStudy (angela210793):

i thought sqr(x^2-9)=t---->x=sqr(t^2+9) dx=(sqrt^2+9)'dt=t/(sqr(t^2+9)dt... won't work???? :S

OpenStudy (angela210793):

http://www.wolframalpha.com/input/?i=integral+1%2F%28%28x^2-9%29^%281%2F2%29%29+dx this is wht Wolfram. says :/

OpenStudy (anonymous):

you guys are unbelievable. where I been all these days?

OpenStudy (anonymous):

But this online stuff says like this, http://www.numberempire.com/definiteintegralcalculator.php

OpenStudy (dumbcow):

the trick for integrating sec(x) is multiply by (sec(x)+tan(x))/(sec(x)+tan(x)) let u = sec(x) +tan(x)

OpenStudy (anonymous):

But assigning x=3 sec@ confuses how boundary limits changes. it doesnt give any answer.

OpenStudy (dumbcow):

hmm, you have to use trig substitution though you have to replace @ with x at the end to evaluate @ = sec^-1(x/3) it will be a little messy

OpenStudy (dumbcow):

after substituting in i get: \[\ln (\frac{\sqrt{x^{2}-9}}{3}+\frac{x}{3})\]

OpenStudy (anonymous):

dumbcow, I tried as you said, I am getting boundary condition pi/6 to pi/4. and answer of integral is ln(x). so once boundary condition applied, then answer is 0.405. am not sure this is right.

OpenStudy (dumbcow):

oh i didn't change the boundary condition since i converted everything back to x

OpenStudy (dumbcow):

After evaluating from 3sqrt2 to 2sqrt3 = 0.332

OpenStudy (anonymous):

you are exactly right. great dumbcow!!

OpenStudy (dumbcow):

are you able to get what i got.. ln(u) = ln(sec(@) + tan(@)) = ln(sec@ + sqrt[sec^2@ - 1]) @ = sec^-1(x/3)

OpenStudy (anonymous):

sorry guys, I slept inbetween., Oops

OpenStudy (anonymous):

thanks for your great ideas.